# Aligning First, Then Fusing: A Novel Weakly Supervised Multimodal Violence Detection Method

### v1

@@ 1 Introduction
The core challenge of weakly supervised multimodal VD lies in effectively integrating information from different modalities in the absence of detailed label information, which can also be framed as the challenge of achieving effective modality fusion. One straightforward fusion approach is to directly concatenate visual and audio features without any processing, as used in earlier multimodal VD methods [9, 10]. However, this approach has clear limitations. These limitations stem from the significant differences between the two modalities, which can be observed in two key aspects: 1) Modality information discrepancy: Visual content typically conveys much more information than audio, and much of the visual content is unrelated to the audio. For example, while a slap might produce a "smack" sound, other details, such as the clothing style and color of individuals, cannot be captured by audio. Treating both modalities equally, without distinguishing their unique contributions, can introduce redundant information and ultimately diminish the value of the audio data. 2) Modality asynchrony: Visual and audio information are sometimes temporally misaligned. For instance, in a typical "abuse" scenario, the abuser strikes the victim first, followed by a scream. Ignoring this temporal misalignment disrupts the causal relationship between the modalities, hindering accurate event detection.
@@
The alignment mentioned above refers to enhancing the consistency of event expression across the two modalities. This involves two aspects: 1) reducing the distance between the semantic features of the two modalities after matching, and 2) ensuring temporal consistency between the modalities. This approach strengthens the modality features that are strongly related to event expression, helping to reduce the impact of redundant information in each modality and improving the complementarity of the modality-specific information. Our complete model further refines the visual modality into RGB video and optical flow video (referred to as RGB and flow). In the context mentioned earlier, "visual" refers to RGB. Similar to the relationship between audio and RGB, flow focuses on dynamic details, whereas RGB contains more comprehensive information. Therefore, its semantic features also exhibit sparse projection in the RGB feature space. Consequently, the complete model searches for two MFMSs: Audio-RGB and Flow-RGB. The alignment process involves pairwise alignment among the three modalities—audio, RGB, and flow—which will be explained in detail in Section 3. As can be seen, our method essentially establishes a primary modality—RGB. The primary modality constructs the main information framework, with other modalities serving as supplementary within this framework, which aligns with the inherent characteristics of each modality.
@@ 2.2 Multimodal Alignment and Fusion
The literature [20] classifies multimodal alignment into two types: Explicit Alignment and Implicit Alignment. A key characteristic of Explicit Alignment is the direct measurement of similarity, while Implicit Alignment generally does not explicitly align the data; instead, alignment is a part of the model’s latent structure in tasks involving modality interactions, such as Image Captioning [27], Visual Question Answering [28], etc. From this perspective, our alignment method is closer to Explicit Alignment. The foundational work in Explicit Alignment includes Canonical Correlation Analysis (CCA) [29] and its deep learning extension, Deep CCA [30]. CCA finds a linear transformation that projects data from two modalities into a shared space and maximizes their correlation, while Deep CCA uses deep neural networks for nonlinear mapping, allowing for better alignment of features across modalities. In recent multimodal learning works, CCA is often used as a loss function to align modalities [31, 32, 33]. Our approach differs from previous methods. The similarity measure aims to identify subspaces in the feature space of the primary modality (the main source of information) that correspond to the secondary modality (which provides complementary information) in the event representation. These subspaces are a part of the primary modality’s feature space and are used to sparsely map the features of the secondary modality into the primary modality’s feature space. The alignment loss used for training (see Section 3.2) primarily considers the features of the primary modality and the sparsified secondary modality, their pairwise distances, and the consistency of the mapping of modality features in the temporal decision sequence space.
@@ 3.1 Unimodal Semantic Feature Extraction
To ensure that these high-level features are strongly correlated with the VD task, we apply a regression layer and MIL loss to further refine them. The regression layer consists of a three-layer MLP, which produces the anomaly score for each time step:
@@
The MIL loss uses the widely adopted top-K strategy: it averages the top-K anomaly scores, i.e., $\bar{s}=\frac{1}{K}\sum_{i\in T_{K}(s)}s_{i}$, where $T_{K}(s)$ represents the set of top-K scores in $s$. Thus, the MIL loss for each modality can be described as follows:
@@ 3.2 Multimodal Alignment
After obtaining the sparse audio and flow features, the first alignment can be performed. This alignment process is achieved by minimizing three losses, which are designed to reduce the pairwise similarity between the three modalities:
@@
Fig. 3: By searching for NFMSs, the entire RGB feature space is divided into four distinct parts: RGB-Audio-Flow MFMS, RGB-Audio MFMS, RGB-Flow MFMS, and pure RGB.
Algorithm 1 Searching for NFMS and sparsifying secondary modality features based on NFMS.
@@
5:Define $\hat{z_{s}}\in\mathbb{R}^{(b\times t)\times d_{s}},\quad\hat{z_{p}}\in\mathbb{R}^{(b\times t)\times d_{p}}.$                                                    # Flatten the inputs into 2D
@@
11:     $S_{\text{top-k}}[i,:]=\text{argsort}(S[i,:])[:k]$       # Where, $\text{argsort}(S[i,:])$ returns the indices of the sorted elements of $S[i,:]$, we set $k=d_{p}$.
@@
21:Define $\hat{\theta}=\{i\in\{1,2,\dots,d_{p}\}\mid i\notin\theta\}$.     # $\hat{\theta}$ as the set of feature dimension indices from $1$ to $d_{p}$ that are not in $\theta$
22:Define $\theta^{\text{pad}}=[\theta,\hat{\theta}]$.                                                                                    # Concatenating $\theta$ and $\hat{\theta}$
@@
$\mathcal{L}_{RA}^{S-\text{CE}}=\text{Score-CrossEntropy}(s_{R},\hat{s_{A}}),$ (22)
$\mathcal{L}_{RF}^{S-\text{CE}}=\text{Score-CrossEntropy}(s_{R},\hat{s_{F}}),$ (23)
$\mathcal{L}_{AF}^{S-\text{CE}}=\text{Score-CrossEntropy}(\hat{s_{A}},\hat{s_{F}}).$ (24)
@@ 3.3 Multimodal Fusion and Violence Detection
The process of computing the triplet loss can be simply expressed as:
$\mathcal{L}_{RAF}^{Triplet}=\Psi(\hat{z_{RAF}},s_{RAF},y),$ (35)
where $\Psi$ represents the process of calculating the triplet loss.
@@ 4.1 Dataset and Evaluation Metric
The XD-Violence [9] dataset is the only large-scale publicly available multimodal dataset for VD to date, containing 4,754 untrimmed videos with a total duration of 217 hours. The training set consists of 3,954 videos, including 1,905 violent and 2,049 non-violent videos, while the test set includes 800 videos, with 500 violent and 300 non-violent videos. The videos in XD-Violence come from various scenes, including movies and YouTube, and cover a range of common violence types, giving the dataset significant advantages in terms of diversity and realism. Each video provides three modalities: RGB video, audio, and flow, offering rich resources for multimodal VD. During inference, we follow prior works [10, 14, 38, 12] and use frame-level AP as the evaluation metric.
@@ 4.2 Implementation Details
4.3 Experimental Results on XD-Violence
@@ 4.5 Qualitative Analysis
(a) Video with riot content
(b) Video with explosion content
(c) Video with fighting
(d) Non-violent street video
(e) Non-violent indoor activity
(f) Normal basketball game
Fig. 7: Anomaly scores of our method on the XD-Violence test videos. Pink areas indicate the manually labelled violence events.
Fig. 7 presents the anomaly detection scores of our method applied to the XD-Violence test videos. The figure includes six different video clip examples, labeled from (a) to (f), to illustrate the detection performance of our method.
Specifically, part (a) shows a video clip containing riot content, where the sections with lower anomaly scores correspond to transitions back to the studio. In the third riot segment of the video, a sudden drop in the anomaly score occurs due to the transition to a relatively calm scene. Part (b) shows a video with an explosion event, where both explosion scenes are correctly identified. Each explosion scene includes both a regular-speed segment and a slow-motion segment. In the slow-motion segment of the first explosion scene, the anomaly score briefly decreases due to the transition to a non-explosive scene, which lacks the corresponding explosion sounds to aid in detection. Part (c) displays a video of a fight during an ice hockey match, with both fighting segments correctly identified. During the match, occasional spikes in the anomaly score are observed, which correspond to intense scenes such as scrambles for the puck in front of the goal. In these cases, the model controls the anomaly score based on audio and other features, keeping the scores low. Part (d) shows a non-violent street activity video, where despite the crowd and noise, the model outputs anomaly scores close to 0. Part (e) demonstrates a non-violent indoor activity video, where the anomaly score only shows small fluctuations during the segments with intense activity. Part (f) presents a normal basketball game video, where the anomaly score only shows minor fluctuations during some intense moments of confrontation.

### v2

@@ 1 Introduction
The core challenge of weakly supervised multimodal VD lies in effectively integrating information from different modalities in the absence of detailed label information, which can also be framed as the challenge of achieving effective modality fusion. One straightforward fusion approach is to directly concatenate visual and audio features without any processing, as used in earlier multimodal VD methods [9, 10]. However, this approach has clear limitations due to the stark differences between these modalities, which manifest themselves in two key aspects: modality information discrepancy and modality asynchrony. These two phenomena are particularly prevalent in VD tasks. In many violent events, the audio modality typically captures transient sounds like hits, gunshots, explosions, or screams, while the visual modality conveys richer, more detailed information, such as color variations, facial expressions, and physical interactions. Moreover, the timing of the modality features may differ; for instance, an attack action might precede the victim’s scream, or in a shooting incident, audio features might occur earlier than the corresponding visual features, even though both convey the same semantic meaning. Ignoring these modality-specific characteristics and treating both modalities equally can lead to redundant information inclusion and ultimately reduce the effectiveness of the audio modality, even misrepresenting the causal relationships across modalities.
@@
The alignment mentioned above refers to enhancing the consistency of event expression across the two modalities. This involves two aspects: 1) reducing the distance between the semantic features of the two modalities after matching, and 2) ensuring temporal consistency between the modalities. This approach strengthens the modality features that are strongly related to event expression, helping to reduce the impact of redundant information in each modality and improving the complementarity of the modality-specific information. Our complete model further refines the visual modality into RGB video and optical flow video (referred to as RGB and flow). In the context mentioned earlier, "visual" refers to RGB. Similar to the relationship between audio and RGB, flow focuses on dynamic details, whereas RGB contains more comprehensive information. Therefore, its semantic features also exhibit sparse projection in the RGB feature space. Consequently, the complete model searches for two MFMSs: RGB-Audio and RGB-Flow. The alignment process involves pairwise alignment among the three modalities—audio, RGB, and flow—which will be explained in detail in Section 3. As can be seen, our method essentially establishes a primary modality—RGB. The primary modality constructs the main information framework, with other modalities serving as supplementary within this framework, which aligns with the inherent characteristics of each modality.
@@ 2.2 Multimodal Alignment and Fusion
The literature [20] classifies multimodal alignment into two types: Explicit Alignment and Implicit Alignment. A key characteristic of Explicit Alignment is the direct measurement of similarity, while Implicit Alignment generally does not explicitly align the data; instead, alignment is a part of the model’s latent structure in tasks involving modality interactions, such as Image Captioning [27], Visual Question Answering [28], etc. From this perspective, our alignment method is closer to Explicit Alignment. The foundational work in Explicit Alignment includes Canonical Correlation Analysis (CCA) [29] and its deep learning extension, Deep CCA [30]. CCA finds a linear transformation that projects data from two modalities into a shared space and maximizes their correlation, while Deep CCA uses deep neural networks for nonlinear mapping, allowing for better alignment of features across modalities. In recent multimodal learning works, CCA is often used as a loss function to align modalities [31, 32, 33]. Our approach differs from previous methods. The similarity measure aims to identify subspaces in the feature space of the primary modality (the main source of information) that correspond to the secondary modality (which provides complementary information) in the event representation. These subspaces are a part of the primary modality’s feature space and are used to sparsely map the features of the secondary modality into the primary modality’s feature space.The alignment loss used for training (see Section 3.2) primarily considers the features of the primary modality and sparsified the secondary modality, their pairwise distances, and the consistency of the mapping of modality features in the temporal decision sequence space.
@@ 3.1 Unimodal Semantic Feature Extraction
To ensure that the high-level features are strongly correlated with the VD task, we treat the training of each modality’s encoder as a weakly supervised VD task. Specifically, we first apply a regression layer at the end of each modality’s encoder to obtain frame-level anomaly scores. This regression layer consists of a three-layer MLP, which computes the anomaly score for each time step:
@@
Due to the lack of frame-level annotations, we adopt a MIL loss, which employs the widely-used top-K strategy: it averages the top-K anomaly scores, i.e., $\bar{s}=\frac{1}{K}\sum_{i\in T_{K}(s)}s_{i}$, where $T_{K}(s)$ represents the set of top-K scores in $s$. Thus, the MIL loss for each modality can be described as follows:
@@ 3.2 Multimodal Alignment
After obtaining the sparse audio and flow features, the first alignment can be performed. This alignment process is achieved by minimizing three losses, which are designed to increase the pairwise similarity between the three modalities:
@@
Fig. 3: By searching for MFMSs, the entire RGB feature space is divided into four distinct parts: RGB-Audio-Flow MFMS, RGB-Audio MFMS, RGB-Flow MFMS, and pure RGB.
Algorithm 1 Searching for MFMS and sparsifying secondary modality features based on MFMS.
@@
5:Define $\hat{z_{s}}\in\mathbb{R}^{(b\times t)\times d_{s}},\quad\hat{z_{p}}\in\mathbb{R}^{(b\times t)\times d_{p}}.$       # Flatten the inputs into 2D.
@@
11:     $S_{\text{top-k}}[i,:]=\text{argsort}(S[i,:])[:k]$       # Where, $\text{argsort}(S[i,:])$ returns the indices of the sorted elements of $S[i,:]$, with $d_{s}\leq k\leq d_{p}$.
@@
21:Define $\hat{\theta}=\{i\in\{1,2,\dots,d_{p}\}\mid i\notin\theta\}$.    # $\hat{\theta}$ as the set of feature dimension indices from $1$ to $d_{p}$ that are not in $\theta$.
22:Define $\theta^{\text{pad}}=[\theta,\hat{\theta}]$.                            # Concatenating $\theta$ and $\hat{\theta}$.
@@
[…]
@@
Eq. (22), (23), and (24) define the loss function that minimizes discrepancies between anomaly score sequences from different modalities, enabling the projection model to capture cross-modal causal relationships. In simpler terms, it identifies features susceptible to modality asynchrony and applies adaptive "compensation" to mitigate inconsistencies.
@@ 3.3 Multimodal Fusion and Violence Detection
The process of computing the triplet loss can be simply expressed as:
$\mathcal{L}_{RAF}^{Triplet}=\Psi(\hat{z_{RAF}},s_{RAF},y),$ (35)
where $\Psi$ represents the process of calculating the triplet loss.
@@ 4.1 Dataset and Evaluation Metric
The XD-Violence [9] dataset is the only large-scale publicly available multimodal dataset for VD to date, containing 4,754 untrimmed videos with a total duration of 217 hours. The training set consists of 3,954 videos, including 1,905 violent and 2,049 non-violent videos, while the test set includes 800 videos, with 500 violent and 300 non-violent videos. The videos in XD-Violence come from various scenes, including movies and YouTube, and cover a range of common violence types, giving the dataset significant advantages in terms of diversity and realism. Each video provides three modalities: RGB video, audio, and flow, offering rich resources for multimodal VD.
Since the XD-Violence test set provides frame-wise annotation labels for the videos, we follow prior works [10, 14, 38, 12] and use frame-level average precision (AP) as the evaluation metric during inference.
@@ 4.2 Implementation Details
4.3 Comparisons with State-of-the-Arts on XD-Violence Dateset
@@
4.4 Training Generalization, Robustness, and Inference Efficiency
Table 2: Performance Comparison on Different Training Set Proportions and Test Sets
Method 30% Training Set 70% Training Set 100% Training Set
Test A Test B Test A Test B Test A Test B
MACIL-SD [14] 80.54 77.32 82.88 79.95 84.73 81.87
HyperVD [44] 80.39 76.42 82.17 77.94 86.93 84.07
Wu et al. [10] 79.76 74.07 80.15 75.95 82.21 76.69
MSBT [12] 82.52 77.63 81.07 78.79 85.48 82.97
Ours 83.56 82.80 85.09 83.03 87.28 84.52
Note: All values are AP (%). Test A/B denote different test subsets.
To evaluate the generalization ability and robustness of our approach, we designed a series of comparative experiments to investigate the accuracy performance of different methods under various training set ratios and testing subsets. Additionally, we analyzed the performance variations of each method when applying different frame drop rates during the training phase. The comparative methods include two typical tri-modal fusion strategies: the direct concatenation fusion method Wu et al. [10] and the weighted fusion method MSBT [12], as well as a unique dual-modal fusion method based on self-distillation, namely MACIL-SD [14].
As shown in Table 2, we randomly selected different proportions (30% and 70%) of samples from the XD-Violence training set to form new training sets, and equally divided the testing set into two subsets, Test A and Test B. Based on this configuration, we compared the accuracy performance of various methods under different training and testing conditions to validate their generalization ability across different data distributions. Experimental results demonstrate that the performance of weakly supervised multimodal violence detection methods is positively correlated with the scale of the training dataset; as the amount of training data increases, the performance of all methods improves. Under the 30% training set condition, our method achieves APs of 83.56% and 82.80% on Test A and Test B, respectively, significantly outperforming the competing methods. When the training set is increased to 70%, our method further raises the AP to 85.09% (Test A) and 83.03% (Test B), which remains markedly higher than that of other methods. This fully demonstrates the superior generalization ability of our approach across different data distributions.
Table 3: Performance Comparison under Varying Training-phase Frame Drop Rates
Method No Drop 10% Drop 30% Drop 50% Drop
MACIL-SD [14] 83.40 80.47 78.46 77.57
HyperVD [44] 85.67 78.67 78.42 76.99
Wu et al. [10] 79.53 78.41 78.57 77.09
MSBT [12] 84.32 79.42 80.06 74.96
Ours 86.07 84.30 81.56 78.22
Note: All values are AP in percentage. Frame drop rates are applied only during training phase, while evaluation uses complete test samples.
We applied different frame drop rates (10%, 30%, and 50%) to the modal data of each sample in the XD-Violence training set to construct new datasets that simulate modality dropout in real-world scenarios, and then observed the performance variations of different methods. As shown in Table 3, although the AP of all methods decreases with increasing frame drop rates, our method consistently outperforms the alternatives. Notably, under 10% and 30% frame drop conditions, our method achieves APs of 84.30% and 81.56%, respectively (while other methods yield APs ranging from 78.41% to 80.47%). Even under the extreme condition of a 50% frame drop rate, our method maintains an AP of 78.22%, surpassing the performance of the other methods. This fully verifies the robustness of our approach in handling incomplete inputs.
Table 4: Comparative Analysis of Model Complexity and Inference Efficiency Across Three Typical Three-Modal Fusion Methods on XD-Violence Benchmark
Method Params (M) Inf. Time (s/sample) AP (%)
Wu et al. [10] 1.43 0.05 79.53
MSBT [12] 15.7 0.19 84.32
Ours 1.36 0.07 86.07
Table 4 presents a comparative analysis of model complexity and inference efficiency for three typical three-modal fusion methods on the XD-Violence benchmark. The table reports the number of parameters (in millions), the average inference time per sample (in seconds), and the detection accuracy (AP%). It can be seen that the direct concatenation fusion method by Wu et al. benefits from its simple structure, achieving the fastest inference speed and a relatively low parameter count, though its detection performance is limited. In contrast, the MSBT method, with its complex modal weighted fusion mechanism, effectively enhances detection accuracy—albeit at the cost of increased model complexity and slower inference. By comparison, our method achieves the highest detection accuracy while maintaining a low overall parameter count and fast inference speed.
@@ 4.5 Ablation Study
Table 6: Ablation Study of Hyperparameters
Index $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ AP (%)
1 1 1 0.001 85.73
2 0.01 85.82
3 1 5 0.001 85.54
4 0.01 85.71
5 1 10 0.001 85.42
6 0.01 85.46
7 10 1 0.001 86.06
8 0.01 86.00
9 10 5 0.001 85.99
10 0.01 85.97
11 10 10 0.001 86.07
12 0.01 85.92
Table 6 presents the results of the ablation study on hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, where the performance is measured by AP (%). $\lambda_{1}$ is the coefficient of $\mathcal{L}_{RAF}^{MIL}$, which is strongly correlated with the final output. Since it plays a crucial role in optimizing the primary objective, we examine whether setting it to 1 or 10 affects the results. $\lambda_{2}$ is the coefficient of $\mathcal{L}_{RAF}^{MA}$, which is closely related to the alignment function. To explore the impact of alignment weight on performance, we consider values of 1, 5, and 10. $\lambda_{3}$ is the coefficient of $\mathcal{L}_{RAF}^{Triplet}$, which mainly enhances the model’s ability to distinguish violent content from non-violent content within violent samples. As it serves as an auxiliary term, we assign it a relatively small weight of 0.001 or 0.01. The results indicate that different combinations of $\lambda_{1}$ and $\lambda_{2}$ lead to slight variations in AP. The highest AP (86.07%) is achieved when $\lambda_{1}=10$, $\lambda_{2}=10$, and $\lambda_{3}=0.001$, suggesting that this setting is optimal among the tested configurations.
@@ 4.6 Qualitative Analysis
Fig. 7: Visualization of the convergence process for RGB-Audio and RGB-Flow MFMS. In (a), the plot shows the evolution of the MFMS convergence indicator during iterations, with the red curve representing RGB-Audio and the green curve representing RGB-Flow. In (b), red bars indicate the frequency with which each RGB feature dimension is selected as part of the RGB-Audio MFMS over 50 iterations—the darker the red, the more frequently it is chosen. Similarly, in (c), green bars depict the frequency with which each RGB feature dimension is identified as part of the RGB-Flow MFMS over 50 iterations, with darker green corresponding to higher frequency.
Fig. 7 visualizes the convergence process for the RGB-Audio and RGB-Flow MFMS. In (a), the plot illustrates the evolution of the MFMS convergence indicator over iterations. The MFMS convergence indicator at the $i$th iteration are computed as follows:
First, a sliding window of length $w$ is defined. The set of MFMS results within this window is given by $\{M_{i-w},M_{i-(w-1)},\ldots,M_{i}\},$ which represents MFMSs identified over the last $w$ iterations starting from iteration $i$.
Let $d_{p}$ denote the dimensionality of the primary modality’s feature space. Within the sliding window, the number of times the $k$th feature (with $1\leq k\leq d_{p}$) of the primary modality is recognized as part of the MFMS is computed as
$f^{k}_{i}=\sum_{j=i-w}^{i}\mathbf{1}_{\{k\in M_{j}\}},$ (37)
where $\mathbf{1}_{\{k\in M_{j}\}}$ is the indicator function that equals 1 if $k\in M_{j}$ and 0 otherwise.
Using these counts, we determine the maximum count $f^{\max}_{i}$ among all features and count the number of features $n_{i}$ that reach this maximum:
$f^{\max}_{i}$ $=\max_{1\leq k\leq d_{p}}f^{k}_{i},$ (38)
$n_{i}$ $=\sum_{k=1}^{d_{p}}\delta(f^{k}_{i}-f^{\max}_{i}),$ (39)
with $\delta(x)$ denoting the Kronecker delta function.
The MFMS convergence indicator capture two aspects: if $f^{\max}_{i}$ is closer to the window length $w$, an MFMS is more likely to exist; and if $n_{i}$ is closer to the secondary modality’s feature dimension $d_{s}$, the distribution of the MFMS is more stable. Therefore, the MFMS convergence indicator at the $i$th iteration is computed by
$\text{m}_{i}=\frac{w\cdot d_{s}}{f^{\max}_{i}\cdot n_{i}}.$ (40)
In (a), the red and green curves represent the convergence indicator for the RGB-Audio and RGB-Flow MFMS, respectively, clearly indicating that the RGB-Flow MFMS converges faster than the RGB-Audio MFMS. (b) and (c) show the changes in the distribution of the RGB-Audio and RGB-Flow MFMS in the RGB feature space during training, further demonstrating that the RGB-Flow MFMS converges more rapidly. This is because the Flow modality can be derived directly from the RGB modality, making it easier to identify the MFMS in the RGB space, whereas the Audio and RGB modalities originate from different sources and have significantly different data structures, resulting in a slower process for identifying the RGB-Audio MFMS in the RGB space.
(i) Refer to caption
(ii) Refer to caption
(d) Video with riot content
(i) Refer to caption
(ii) Refer to caption
(e) Video with explosion
(i) Refer to caption
(ii) Refer to caption
(f) Video with fighting
Fig. 8: Anomaly scores of our method on XD-Violence test videos with violent content. The figure shows three test videos, each with two parts: (i) A comparison between the model’s final anomaly scores output and the ground truth, with pink-highlighted regions indicating manually annotated violence events. (ii) A comparison of anomaly scores for individual modalities and the fused anomaly scores after modality fusion.
Fig. 8 presents the anomaly score curves of our method on three violent videos from the XD-Violence test set. Specifically, (a) shows riot scenes from news reports, (b) depicts explosions in TV shows, and (c) illustrates fights in ice hockey matches. Each case consists of two subplots: (i) a comparison between the model’s final anomaly scores and the ground truth labels, with pink-highlighted regions indicating the annotated violent segments; and (ii) a comparison of the anomaly score curves produced by each modality’s regression layer against the final anomaly score curve after modality fusion, where blue represents the audio modality, gold the flow modality, green the RGB modality, and red the output of the fused model.
In the riot scene video (a), the model successfully identifies the violent segments, with one exception: the third riot scene, where the anomaly score sharply decreases due to a transition to a relatively calm scene. Further analysis of subplot (ii) reveals that the audio modality is highly responsive to abnormal sounds like explosions and gunshots, while the flow modality effectively detects all riot segments but shows a weak response to the low-intensity third scene. The RGB modality, on the other hand, experiences a delay when transitioning from violent to non-violent segments. However, the fusion model effectively leverages the strengths of each modality, accurately distinguishing between violent and non-violent segments and overcoming the individual limitations of each modality.
In the explosion scene (b), the model successfully detects both regular-speed and slow-motion explosions. However, similar to video (a), the anomaly score sharply drops for the first slow-motion explosion segment due to a sudden camera switch to a non-explosive scene. Analysis of subplot (ii) shows varying performance across modalities: the audio modality struggles with slow-motion explosions due to low-frequency noise masking the acoustic features but remains highly responsive to regular-speed explosions; the flow modality, benefiting from enhanced temporal continuity in slow-motion segments, accurately captures slow-motion explosions but misses regular-speed ones; and the RGB modality experiences two false positives due to camera switching. Nevertheless, the fusion model effectively combines the strengths of each modality, accurately identifying both regular-speed and slow-motion explosions while minimizing false positives.
In the ice hockey match scene (c), the model accurately identifies the two fight segments. Analysis of subplot (ii) reveals that the audio modality’s baseline anomaly score is biased high due to continuous impact sounds, resulting in many false positives. The flow modality shows minimal response during violent periods but remains stable and low during non-violent segments. The RGB modality experiences periodic noise from rapid camera switching but responds accurately to both violent segments. The fusion model, by leveraging the strengths of all modalities, effectively filters out noise, accurately identifies the violent segments, and reduces false positives.
(a) Documentary clip
(b) Indoor dance
(c) Crowd movement
Fig. 9: Anomaly score distribution on non-violent XD-Violence test scenarios. Three representative videos are shown, each containing comparative visualization of unimodal anomaly scores and fused scores after modality fusion.
Fig. 9 shows the anomaly score curves of our method on the XD-Violence non-violent test scenes. By analyzing three typical cases—documentary footage (a), indoor dance (b), and crowd movement (c)—we compare the outputs of the unimodal regression layers with the multimodal fusion results. Blue, gold, and green represent the audio, flow, and RGB modalities, respectively, while red indicates the fusion model’s outputs.
In the documentary footage (a), the audio modality remains stable. The flow modality shows a slight increase in the anomaly score during the strobe light effect used by the creator, while the RGB modality experiences a sudden spike in anomaly score due to rapid camera changes. The fusion model effectively suppresses these noises, with the largest fluctuation in anomaly scores staying under 0.2.
In the indoor dance scene (b), the audio modality’s regression layer misinterprets the frequent sounds of high heels hitting the floor, causing an abnormal increase in the anomaly score. Neither the flow nor the RGB modality responds to this. However, the fusion model successfully suppresses this misjudgment, resulting in a stable output.
In the crowd movement scene (c), both the audio and flow modalities remain stable, while the RGB modality slightly misjudges the intense motion caused by a door opening. The fusion model effectively suppresses this error, providing a more stable output.
The experimental findings from both violent and non-violent scenarios demonstrate that, after aligning the features of each modality, the fusion model effectively harnesses their strengths, thereby enhancing the model’s robustness in various complex scenarios.